OpenStudy (anonymous):

What is the slope and y-intercept of the graph of -x + 2y = 6?

OpenStudy (anonymous):

I am confused can someone please help

OpenStudy (aripotta):

convert the equation to slope-intercept form, and you should be able to easily find the answers

OpenStudy (aripotta):

slope-intercept is y = mx + b, where m is the slope and b is the y-intercept.

OpenStudy (anonymous):

would it be 1, 3

OpenStudy (aripotta):

-x + 2y = 6 add x to both sides: 2y = x + 6 divide both sides by 2: y = 1/2x + 3

OpenStudy (aripotta):

so the answer is...?

OpenStudy (anonymous):

1/2, 3

OpenStudy (aripotta):

yep. do you see how i converted the equation?